OpenStudy (anonymous):

Darla is building a deak. To make sure she had made a square corner, she measures 9 ft. from the corner along one edge & 8ft from the corner along the other edge. How long should the diagonal between those two points if the corner is a right angle?

OpenStudy (anonymous):

*desk

OpenStudy (anonymous):

In a right angle triangle, Pythagoras told us that the hypotenuse is the root of the sum of the squares of the other two sides. a^2+b^2=c^2. You have a = 9ft (108") and b = 8ft (96") square them, sum them and take the root to get c.